Find the values of a for which the function f (x) = cos x - sin x + ax + b is decreasing

Grade:12

2 Answers

Sher Mohammad IIT Delhi
9 years ago
f'(x)=-sinx-cosx+a <0
the minmum value of -sinx-cosx, occurs at pi/4 = -sqrt(2)=-1.41
a-1.41<0
a<1.41

Sher Mohammad
B.Tech, IIT Delhi
Sher Mohammad IIT Delhi
9 years ago
'(x)=-sinx-cox+a
minimum of -sinx-cosx= -sqrt(2)=-1.4


f'(x)>a-1.4
f'(x)>0
a-1.4>=0
a belongs to(1.4, inf)

Sher Mohammad
B.tech, IIT Delhi